Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 06:02:32 on July 06, 2010
Entry number 330395
keyword=Owl shift summary
Owl Shift Summary: July 6, 2010
- Pete Markowitz
- James Beacham
- Victor Maxwell
- Lawrence Selvy
- 00:00 Hall is in controlled access to reset RHRS Q1.
- 00:26 Septum back on, ask for beam, change to Pb-D1. Raster 6x6,
1uA current. Start run 1784, check spot. Run 1785 is junk
- 00:44 Decrease raster to 4x4 and moved to 0,1 in BPMA and BPMB.
Start run 1786.
- 01:35 The electron-positron tof peak (run 1786) appears to have
worse S/N on the lead than what we saw on the Ta (in run 1780). Natalia
estimates we get about 3Hz in the peak for the lead, vs ~10 Hz for the Ta
(my estimate). [See entry 330373 for the lead, and 330358 for the
tantalum.] We are going back to Ta.
- 01:51 Start run 1787. RHRS Q1 trips off again. Access to reset.
Run is possibly junk.
- 02:10 Turn on septum, ask for beam back. Start run 1788.
- 03:25 Start run 1789.
- 04:50 Start run 1790
- 06:00 MCC takes beam for beam studies.